# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v4

@@ Appendix E Examples output of the planning phase
Paper
Source Score
Generative Judge for Evaluating Alignment
Poster 4
Distributional Preference Learning: Understanding and Accounting for Hidden Context in RLHF
Poster 4
Inherently Interpretable Time Series Classification via Multiple Instance Learning
Oral 3.9
iTransformer: Inverted Transformers Are Effective for Time Series Forecasting
Oral 3.9
Tell Your Model Where to Attend: Post-hoc Attention Steering for LLMs
Poster 3.9
Knowledge Distillation Based on Transformed Teacher Matching
Poster 3.9
Meaning Representations from Trajectories in Autoregressive Models
Poster 3.8
A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Poster 3.8
VDC: Versatile Data Cleanser based on Visual-Linguistic Inconsistency by Multimodal Large Language Models
Poster 3.8
Vocos: Closing the gap between time-domain and Fourier-based neural vocoders for high-quality audio synthesis
Poster 3.8
SliceGPT: Compress Large Language Models by Deleting Rows and Columns
Poster 3.8
Beyond Accuracy: Evaluating Self-Consistency of Code Large Language Models with IdentityChain
Poster 3.8
Guiding Masked Representation Learning to Capture Spatio-Temporal Relationship of Electrocardiogram
Poster 3.8
Social Reward: Evaluating and Enhancing Generative AI through Million-User Feedback from an Online Creative Community
Oral 3.7
Language Model Detectors Are Easily Optimized Against
Poster 3.7
Improving protein optimization with smoothed fitness landscapes
Poster 3.7
SparseFormer: Sparse Visual Recognition via Limited Latent Tokens
Poster 3.7
AutoVP: An Automated Visual Prompting Framework and Benchmark
Poster 3.7
Hierarchical Context Merging: Better Long Context Understanding for Pre-trained LLMs
Poster 3.7
SEABO: A Simple Search-Based Method for Offline Imitation Learning
Poster 3.7
OpenChat: Advancing Open-source Language Models with Mixed-Quality Data
Poster 3.7
Rethinking The Uniformity Metric in Self-Supervised Learning
Poster 3.7
VONet: Unsupervised Video Object Learning With Parallel U-Net Attention and Object-wise Sequential VAE
Poster 3.6
Efficient Backpropagation with Variance-Controlled Adaptive Sampling
Poster 3.6
Structuring Representation Geometry with Rotationally Equivariant Contrastive Learning
Poster 3.6
ControlVideo: Training-free Controllable Text-to-Video Generation
Poster 3.6
Context-Aware Meta-Learning
Poster 3.6
RECOMBINER: Robust and Enhanced Compression with Bayesian Implicit Neural Representations
Poster 3.6
Peering Through Preferences: Unraveling Feedback Acquisition for Aligning Large Language Models
Poster 3.6
Modulate Your Spectrum in Self-Supervised Learning
Poster 3.6
@@ A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Paper
Source Score
SAMformer: Unlocking the Potential of Transformers in Time Series Forecasting with Sharpness-Aware Minimization and Channel-Wise Attention
Oral 4
Autoformalizing Euclidean Geometry
Poster 4
Recurrent Distance Filtering for Graph Representation Learning
Poster 4
CosPGD: an efficient white-box adversarial attack for pixel-wise prediction tasks
Poster 3.9
Token-level Direct Preference Optimization
Poster 3.9
BayOTIDE: Bayesian Online Multivariate Time Series Imputation with Functional Decomposition
Oral 3.8
CurBench: Curriculum Learning Benchmark
Poster 3.8
Exploring the Low-Pass Filtering Behavior in Image Super-Resolution
Poster 3.8
Towards Efficient Exact Optimization of Language Model Alignment
Poster 3.7
On the Effectiveness of Supervision in Asymmetric Non-Contrastive Learning
Poster 3.7
Drug Discovery with Dynamic Goal-aware Fragments
Poster 3.7
Fool Your (Vision and) Language Model With Embarrassingly Simple Permutations
Poster 3.7
Image Restoration Through Generalized Ornstein-Uhlenbeck Bridge
Poster 3.7
Timer: Generative Pre-trained Transformers Are Large Time Series Models
Poster 3.7
Mitigating Oversmoothing Through Reverse Process of GNNs for Heterophilic Graphs
Poster 3.7
Scribble-Supervised Semantic Segmentation with Prototype-based Feature Augmentation
Poster 3.7
ConvNet vs Transformer, Supervised vs CLIP: Beyond ImageNet Accuracy
Poster 3.7
CLIF: Complementary Leaky Integrate-and-Fire Neuron for Spiking Neural Networks
Oral 3.6
FiT: Flexible Vision Transformer for Diffusion Model
Oral 3.6
Decomposing Uncertainty for Large Language Models through Input Clarification Ensembling
Oral 3.6
SparseTSF: Modeling Long-term Time Series Forecasting with *1k* Parameters
Oral 3.6
Sample-specific Masks for Visual Reprogramming-based Prompting
Oral 3.6
Boundary Exploration for Bayesian Optimization With Unknown Physical Constraints
Poster 3.6
Listwise Reward Estimation for Offline Preference-based Reinforcement Learning
Poster 3.6
Graph Distillation with Eigenbasis Matching
Poster 3.6
Temporal Spiking Neural Networks with Synaptic Delay for Graph Reasoning
Poster 3.6
Position: Quo Vadis, Unsupervised Time Series Anomaly Detection?
Poster 3.6
Neural SPH: Improved Neural Modeling of Lagrangian Fluid Dynamics
Poster 3.6
Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models
Poster 3.6
Unveiling and Harnessing Hidden Attention Sinks: Enhancing Large Language Models without Training through Attention Calibration
Poster 3.6
@@
Paper
Source Score
PACE: marrying generalization in PArameter-efficient fine-tuning with Consistency rEgularization
Oral 4
The Road Less Scheduled
Oral 4
G-Retriever: Retrieval-Augmented Generation for Textual Graph Understanding and Question Answering
Poster 4
Binarized Diffusion Model for Image Super-Resolution
Poster 4
Learning to Predict Structural Vibrations
Poster 4
Attack-Aware Noise Calibration for Differential Privacy
Poster 4
Make Your LLM Fully Utilize the Context
Poster 3.9
Smoothed Energy Guidance: Guiding Diffusion Models with Reduced Energy Curvature of Attention
Poster 3.9
Sm: enhanced localization in Multiple Instance Learning for medical imaging classification
Poster 3.9
AutoTimes: Autoregressive Time Series Forecasters via Large Language Models
Poster 3.9
End-to-End Ontology Learning with Large Language Models
Poster 3.8
Scaling transformer neural networks for skillful and reliable medium-range weather forecasting
Poster 3.8
Autoregressive Image Generation without Vector Quantization
Oral 3.7
Adaptive Randomized Smoothing: Certified Adversarial Robustness for Multi-Step Defences
Oral 3.7
Generalizable Person Re-identification via Balancing Alignment and Uniformity
Poster 3.7
Universal Neural Functionals
Poster 3.7
Are Self-Attentions Effective for Time Series Forecasting?
Poster 3.7
xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology
Poster 3.7
Leveraging Environment Interaction for Automated PDDL Translation and Planning with Large Language Models
Poster 3.7
Task-Agnostic Machine Learning-Assisted Inference
Poster 3.7
Make Continual Learning Stronger via C-Flat
Poster 3.7
DARG: Dynamic Evaluation of Large Language Models via Adaptive Reasoning Graph
Poster 3.7
AsyncDiff: Parallelizing Diffusion Models by Asynchronous Denoising
Poster 3.7
You Only Look Around: Learning Illumination Invariant Feature for Low-light Object Detection
Poster 3.6
MutaPLM: Protein Language Modeling for Mutation Explanation and Engineering
Poster 3.6
Advancing Training Efficiency of Deep Spiking Neural Networks through Rate-based Backpropagation
Poster 3.6
Improved off-policy training of diffusion samplers
Poster 3.6
Navigating the Effect of Parametrization for Dimensionality Reduction
Poster 3.6
Long-Range Feedback Spiking Network Captures Dynamic and Static Representations of the Visual Cortex under Movie Stimuli
Poster 3.6
InfLLM: Training-Free Long-Context Extrapolation for LLMs with an Efficient Context Memory
Poster 3.6
@@
RepoName
Paper
Score
VideoICL
VideoICL: Confidence-based Iterative In-context Learning for Out-of-Distribution Video Understanding
2.6
MuDI
Identity Decoupling for Multi-Subject Personalization of Text-to-Image Models
3.3
KALMV
Knowledge-Augmented Language Model Verification
3.3
sea-attention
SEA: Sparse Linear Attention with Estimated Attention Mask
2.7
HarmAug
HarmAug: Effective Data Augmentation for Knowledge Distillation of Safety Guard Models
3.0
GruM
Graph Generation with Diffusion Mixture
3.7
Adaptive-RAG
Adaptive-RAG: Learning to Adapt Retrieval-Augmented Large Language Models through Question Complexity
2.7
SoT
Sketch-of-Thought: Efficient LLM Reasoning with Adaptive Cognitive-Inspired Sketching
4.0
Mol-LLaMA
Mol-LLaMA: Towards General Understanding of Molecules in Large Molecular Language Model
3.5
judge_code_efficiency
Rethinking Code Refinement: Learning to Judge Code Efficiency
3.1
KARD
Knowledge-Augmented Reasoning Distillation for Small Language Models in Knowledge-Intensive Tasks
3.2
COINCIDE_code
Concept-skill Transferability-based Data Selection for Large Vision-Language Models
3.0
Janus
Aligning to thousands of preferences via system message generalization
3.5
N/A
Silent Branding Attack: Trigger-free Data Poisoning Attack on Text-to-Image Diffusion Models
N/A
VideoRAG
VideoRAG: Retrieval-Augmented Generation over Video Corpus
3.0
RADA
Retrieval-augmented data augmentation for low-resource domain tasks
3.0
STELLA_code
STELLA: Continual Audio-Video Pre-training with Spatio-Temporal Localized Alignment
3.3
prometheus-vision
Prometheus-vision: Vision-language model as a judge for fine-grained evaluation
3.1
CoLoR
Efficient Long Context Language Model Retrieval with Compression
3.0
Volcano
Volcano: Mitigating Multimodal Hallucination through Self-Feedback Guided Revision
3.2
N/A
T1: Tool-integrated Self-verification for Test-time Compute Scaling in Small Language Models
N/A
@@
Repo Name
Paper
CoLoR
Efficient Long Context Language Model Retrieval with Compression
cognitive-behaviors
Cognitive Behaviors that Enable Self-Improving Reasoners, or, Four Habits of Highly Effective STaRs
RADA
Retrieval-Augmented Data Augmentation for Low-Resource Domain Tasks
Self-Instruct
Self-Instruct: Aligning Language Models with Self-Generated Instructions
G-EVAL
G-Eval: NLG Evaluation using GPT-4 with Better Human Alignment

### v5

@@ Appendix D Prompts
Prompt for addressing the environment in the coding stage [System] You are an expert researcher and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive configuration file named "config.yaml", and implmented code repository. Your task is to write a Bash script that can run the given repository from scratch. The script should create and activate the required environment, install all dependencies, and include the commands needed to execute the main file or entry point. Make sure the script is self-contained and can be executed without any manual setup. Write code with triple quoto. [User] # Context ## Configuration file “‘yaml {The content of the configuration file} “‘ —– ## Code Files {The content of the generated repository} —– # Format example ## Code: reproduce.sh “‘bash ## reproduce.sh “‘ —– # Instruction Based on the code files, follow "Format example", write the code. We have {done_file_lst}. Next, you must write only the "reproduce.sh". ## Code: reproduce.sh
Figure 33: Prompt for addressing the environment in the coding stage. {{}} indicate placeholders to be filled with the content described in the accompanying explanation.
@@ Appendix E Examples output of the planning phase
Paper Source Score
Generative Judge for Evaluating Alignment Poster 4
Distributional Preference Learning: Understanding and Accounting for Hidden Context in RLHF Poster 4
Inherently Interpretable Time Series Classification via Multiple Instance Learning Oral 3.9
iTransformer: Inverted Transformers Are Effective for Time Series Forecasting Oral 3.9
Tell Your Model Where to Attend: Post-hoc Attention Steering for LLMs Poster 3.9
Knowledge Distillation Based on Transformed Teacher Matching Poster 3.9
Meaning Representations from Trajectories in Autoregressive Models Poster 3.8
A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis Poster 3.8
VDC: Versatile Data Cleanser based on Visual-Linguistic Inconsistency by Multimodal Large Language Models Poster 3.8
Vocos: Closing the gap between time-domain and Fourier-based neural vocoders for high-quality audio synthesis Poster 3.8
SliceGPT: Compress Large Language Models by Deleting Rows and Columns Poster 3.8
Beyond Accuracy: Evaluating Self-Consistency of Code Large Language Models with IdentityChain Poster 3.8
Guiding Masked Representation Learning to Capture Spatio-Temporal Relationship of Electrocardiogram Poster 3.8
Social Reward: Evaluating and Enhancing Generative AI through Million-User Feedback from an Online Creative Community Oral 3.7
Language Model Detectors Are Easily Optimized Against Poster 3.7
Improving protein optimization with smoothed fitness landscapes Poster 3.7
SparseFormer: Sparse Visual Recognition via Limited Latent Tokens Poster 3.7
AutoVP: An Automated Visual Prompting Framework and Benchmark Poster 3.7
Hierarchical Context Merging: Better Long Context Understanding for Pre-trained LLMs Poster 3.7
SEABO: A Simple Search-Based Method for Offline Imitation Learning Poster 3.7
OpenChat: Advancing Open-source Language Models with Mixed-Quality Data Poster 3.7
Rethinking The Uniformity Metric in Self-Supervised Learning Poster 3.7
VONet: Unsupervised Video Object Learning With Parallel U-Net Attention and Object-wise Sequential VAE Poster 3.6
Efficient Backpropagation with Variance-Controlled Adaptive Sampling Poster 3.6
Structuring Representation Geometry with Rotationally Equivariant Contrastive Learning Poster 3.6
ControlVideo: Training-free Controllable Text-to-Video Generation Poster 3.6
Context-Aware Meta-Learning Poster 3.6
RECOMBINER: Robust and Enhanced Compression with Bayesian Implicit Neural Representations Poster 3.6
Peering Through Preferences: Unraveling Feedback Acquisition for Aligning Large Language Models Poster 3.6
Modulate Your Spectrum in Self-Supervised Learning Poster 3.6
@@
Paper Source Score
SAMformer: Unlocking the Potential of Transformers in Time Series Forecasting with Sharpness-Aware Minimization and Channel-Wise Attention Oral 4
Autoformalizing Euclidean Geometry Poster 4
Recurrent Distance Filtering for Graph Representation Learning Poster 4
CosPGD: an efficient white-box adversarial attack for pixel-wise prediction tasks Poster 3.9
Token-level Direct Preference Optimization Poster 3.9
BayOTIDE: Bayesian Online Multivariate Time Series Imputation with Functional Decomposition Oral 3.8
CurBench: Curriculum Learning Benchmark Poster 3.8
Exploring the Low-Pass Filtering Behavior in Image Super-Resolution Poster 3.8
Towards Efficient Exact Optimization of Language Model Alignment Poster 3.7
On the Effectiveness of Supervision in Asymmetric Non-Contrastive Learning Poster 3.7
Drug Discovery with Dynamic Goal-aware Fragments Poster 3.7
Fool Your (Vision and) Language Model With Embarrassingly Simple Permutations Poster 3.7
Image Restoration Through Generalized Ornstein-Uhlenbeck Bridge Poster 3.7
Timer: Generative Pre-trained Transformers Are Large Time Series Models Poster 3.7
Mitigating Oversmoothing Through Reverse Process of GNNs for Heterophilic Graphs Poster 3.7
Scribble-Supervised Semantic Segmentation with Prototype-based Feature Augmentation Poster 3.7
ConvNet vs Transformer, Supervised vs CLIP: Beyond ImageNet Accuracy Poster 3.7
CLIF: Complementary Leaky Integrate-and-Fire Neuron for Spiking Neural Networks Oral 3.6
FiT: Flexible Vision Transformer for Diffusion Model Oral 3.6
Decomposing Uncertainty for Large Language Models through Input Clarification Ensembling Oral 3.6
SparseTSF: Modeling Long-term Time Series Forecasting with *1k* Parameters Oral 3.6
Sample-specific Masks for Visual Reprogramming-based Prompting Oral 3.6
Boundary Exploration for Bayesian Optimization With Unknown Physical Constraints Poster 3.6
Listwise Reward Estimation for Offline Preference-based Reinforcement Learning Poster 3.6
Graph Distillation with Eigenbasis Matching Poster 3.6
Temporal Spiking Neural Networks with Synaptic Delay for Graph Reasoning Poster 3.6
Position: Quo Vadis, Unsupervised Time Series Anomaly Detection? Poster 3.6
Neural SPH: Improved Neural Modeling of Lagrangian Fluid Dynamics Poster 3.6
Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models Poster 3.6
Unveiling and Harnessing Hidden Attention Sinks: Enhancing Large Language Models without Training through Attention Calibration Poster 3.6
@@
Paper Source Score
PACE: marrying generalization in PArameter-efficient fine-tuning with Consistency rEgularization Oral 4
The Road Less Scheduled Oral 4
G-Retriever: Retrieval-Augmented Generation for Textual Graph Understanding and Question Answering Poster 4
Binarized Diffusion Model for Image Super-Resolution Poster 4
Learning to Predict Structural Vibrations Poster 4
Attack-Aware Noise Calibration for Differential Privacy Poster 4
Make Your LLM Fully Utilize the Context Poster 3.9
Smoothed Energy Guidance: Guiding Diffusion Models with Reduced Energy Curvature of Attention Poster 3.9
Sm: enhanced localization in Multiple Instance Learning for medical imaging classification Poster 3.9
AutoTimes: Autoregressive Time Series Forecasters via Large Language Models Poster 3.9
End-to-End Ontology Learning with Large Language Models Poster 3.8
Scaling transformer neural networks for skillful and reliable medium-range weather forecasting Poster 3.8
Autoregressive Image Generation without Vector Quantization Oral 3.7
Adaptive Randomized Smoothing: Certified Adversarial Robustness for Multi-Step Defences Oral 3.7
Generalizable Person Re-identification via Balancing Alignment and Uniformity Poster 3.7
Universal Neural Functionals Poster 3.7
Are Self-Attentions Effective for Time Series Forecasting? Poster 3.7
xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology Poster 3.7
Leveraging Environment Interaction for Automated PDDL Translation and Planning with Large Language Models Poster 3.7
Task-Agnostic Machine Learning-Assisted Inference Poster 3.7
Make Continual Learning Stronger via C-Flat Poster 3.7
DARG: Dynamic Evaluation of Large Language Models via Adaptive Reasoning Graph Poster 3.7
AsyncDiff: Parallelizing Diffusion Models by Asynchronous Denoising Poster 3.7
You Only Look Around: Learning Illumination Invariant Feature for Low-light Object Detection Poster 3.6
MutaPLM: Protein Language Modeling for Mutation Explanation and Engineering Poster 3.6
Advancing Training Efficiency of Deep Spiking Neural Networks through Rate-based Backpropagation Poster 3.6
Improved off-policy training of diffusion samplers Poster 3.6
Navigating the Effect of Parametrization for Dimensionality Reduction Poster 3.6
Long-Range Feedback Spiking Network Captures Dynamic and Static Representations of the Visual Cortex under Movie Stimuli Poster 3.6
InfLLM: Training-Free Long-Context Extrapolation for LLMs with an Efficient Context Memory Poster 3.6
@@
RepoName Paper Score
VideoICL VideoICL: Confidence-based Iterative In-context Learning for Out-of-Distribution Video Understanding 2.6
MuDI Identity Decoupling for Multi-Subject Personalization of Text-to-Image Models 3.3
KALMV Knowledge-Augmented Language Model Verification 3.3
sea-attention SEA: Sparse Linear Attention with Estimated Attention Mask 2.7
HarmAug HarmAug: Effective Data Augmentation for Knowledge Distillation of Safety Guard Models 3.0
GruM Graph Generation with Diffusion Mixture 3.7
Adaptive-RAG Adaptive-RAG: Learning to Adapt Retrieval-Augmented Large Language Models through Question Complexity 2.7
SoT Sketch-of-Thought: Efficient LLM Reasoning with Adaptive Cognitive-Inspired Sketching 4.0
Mol-LLaMA Mol-LLaMA: Towards General Understanding of Molecules in Large Molecular Language Model 3.5
judge_code_efficiency Rethinking Code Refinement: Learning to Judge Code Efficiency 3.1
KARD Knowledge-Augmented Reasoning Distillation for Small Language Models in Knowledge-Intensive Tasks 3.2
COINCIDE_code Concept-skill Transferability-based Data Selection for Large Vision-Language Models 3.0
Janus Aligning to thousands of preferences via system message generalization 3.5
N/A Silent Branding Attack: Trigger-free Data Poisoning Attack on Text-to-Image Diffusion Models N/A
VideoRAG VideoRAG: Retrieval-Augmented Generation over Video Corpus 3.0
RADA Retrieval-augmented data augmentation for low-resource domain tasks 3.0
STELLA_code STELLA: Continual Audio-Video Pre-training with Spatio-Temporal Localized Alignment 3.3
prometheus-vision Prometheus-vision: Vision-language model as a judge for fine-grained evaluation 3.1
CoLoR Efficient Long Context Language Model Retrieval with Compression 3.0
Volcano Volcano: Mitigating Multimodal Hallucination through Self-Feedback Guided Revision 3.2
N/A T1: Tool-integrated Self-verification for Test-time Compute Scaling in Small Language Models N/A
@@
Repo Name Paper
CoLoR Efficient Long Context Language Model Retrieval with Compression
cognitive-behaviors Cognitive Behaviors that Enable Self-Improving Reasoners, or, Four Habits of Highly Effective STaRs
RADA Retrieval-Augmented Data Augmentation for Low-Resource Domain Tasks
Self-Instruct Self-Instruct: Aligning Language Models with Self-Generated Instructions
G-EVAL G-Eval: NLG Evaluation using GPT-4 with Better Human Alignment